# Using Large Language Models for Hyperparameter Optimization

### v1

@@ Abstract
This paper studies using foundational large language models (LLMs) to make decisions during hyperparameter optimization (HPO). Empirical evaluations demonstrate that in settings with constrained search budgets, LLMs can perform comparably or better than traditional HPO methods like random search and Bayesian optimization on standard benchmarks. Furthermore, we propose to treat the code specifying our model as a hyperparameter, which the LLM outputs, going beyond the capabilities of existing HPO approaches. Our findings suggest that LLMs are a promising tool for improving efficiency in the traditional decision-making problem of hyperparameter optimization.
@@ 1 Introduction
Hyperparameters, distinct from the parameters directly learned throughout training, significantly influence the inductive bias of a machine learning model, thus determining its capacity to generalize effectively. The choice of hyperparameters helps dictate the complexity of the model, the strength of regularization, the optimization strategy, the loss function, and more. Due to this importance, many methods have been proposed to find strong hyperparameter configurations automatically. For example, black-box optimization methods such as random search (Bergstra and Bengio, 2012) or Bayesian optimization (Mockus, 1998; Shahriari et al., 2015) have been deployed for hyperparameter optimization (HPO). Beyond black-box methods, other works further utilize the structure of the problem, e.g., are compute-environment-aware Ginsbourger et al. (2010), compute-budget-aware (Lam et al., 2016), multi-task (Swersky et al., 2013), like transfer-learning (Golovin et al., 2017) or multi-fidelity (Klein et al., 2017), iterative-optimization-aware (Li et al., 2017), online (Jaderberg et al., 2017), or multi-objective (Daulton et al., 2022). However, these methods (a) still rely on practitioners to design a search space, which includes selecting which parameters can be optimized and specifying bounds on these parameters, and (b) typically struggle in the initial search phase ($<2^{d}$ queries for $d$-dimensional hyperparameters).
@@
This paper investigates the capacity of large language models (LLMs) to optimize hyperparameters. Large language models are trained on internet-scale data and have demonstrated emergent capabilities in new settings (Brown et al., 2020; OpenAI, 2023). We prompt LLMs with an initial set of instructions – describing the specific dataset, model, and hyperparameters – to recommend hyperparameters to evaluate. Upon receiving these hyperparameters, we train the model based on the proposed configuration, record the final metric (e.g., validation loss), and ask the LLM to suggest the next set of hyperparameters. Figure 1 illustrates this iterative process.
@@
Empirically, we employ our proposed algorithm to investigate whether LLMs can optimize 2D toy optimization objectives, where LLMs only receive loss $\mathcal{L}(\boldsymbol{x})$ at specific points $\boldsymbol{x}$. Across a range of objectives, we show that LLMs effectively minimize the loss, exploiting performant regions or exploring untested areas. Next, to analyze whether LLMs can optimize realistic HPO settings, we evaluate our approach on standard HPO benchmarks, comparing it to common methods such as random search and Bayesian optimization. With small search budgets (e.g., $10$ evaluations), LLMs can improve traditional hyperparameter tuning. We also assess the importance of using chain-of-thought reasoning (Wei et al., 2022) and demonstrate that LLMs can also perform well over longer horizons of 30 evaluations. This opens the door to many possibilities, such as using the LLM output to complement other HPO methods.
@@
We further explore the possibilities of optimizing hyperparameters with more flexibility. Rather than adhering to a fixed set of hyperparameter configurations, we prompt LLMs to produce training code (e.g., in PyTorch) to improve validation performance. This extension alleviates the need for human specification of the hyperparameters and their search spaces. As such, code generation can help automate applying machine learning to real-world problems. The generated code provides a strong initialization and can reduce the initial search for configuration spaces that are unlikely to succeed. With a limited search budget ($5$ evaluations), our results show that code generation performs better than random search. We conclude with a discussion of language models as general-purpose hyperparameter tuning assistants, limitations, and future work.
@@ 2 Method
ptLLM AgentptEnvironment Initial prompt: You are helping tune hyperparameters to minimize loss… this is our search space… formulate your output as JSON. Hyperparameter configuration Validation metric e.g., loss
Figure 1: LLMs for hyperparameter optimization. In this sequential decision making setup, we prompt an LLM with the problem description and the search space. The LLM then outputs a set of hyperparameters to evaluate. The environment executes a training run with the hyperparameter setting, and then a validation metric is used to prompt the language model again.
Refer to captionMinimum validation error achieved by neural networks with different optimizers Error RateError Rate
Figure 2: Minimum validation error achieved after $10$ function evaluations from different hyperparameter optimizers on datasets from HPOBench (Eggensperger et al., 2021). The benchmark defines a configuration search space on neural networks for learning rates, $\ell_{2}$ regularization, width, depth, and batch size. A large language model (GPT-4 Turbo) performs well compared to random search and Bayesian optimization ($BO$) with Gaussian process ($GP$) and random forest ($RF$) surrogate models. We also conduct experiments to tune the hyperparameters for logistic regression, support vector machines (SVMs), and random forests, with aggregate results in Table 1.
@@ 2.1 Background
$\boldsymbol{\lambda}^{\star}$ $=\operatorname*{arg\,min}_{\boldsymbol{\lambda}}\mathcal{L}_{V}^{*}\left(\boldsymbol{\lambda}\right)=\operatorname*{arg\,min}_{\boldsymbol{\lambda}}\mathcal{L}_{V}\left(\boldsymbol{\lambda},\mathbf{w}^{\star}\left(\boldsymbol{\lambda}\right)\right)$ (1)
[…]
where $\mathcal{L}_{T}$ and $\mathcal{L}_{V}$ are training and validation objectives, and $\boldsymbol{\lambda}$ and $\mathbf{w}$ are hyperparameters and the model parameters, respectively. Intuitively, the objective – i.e., the validation loss with best-responding parameters $\mathcal{L}_{V}^{*}$ – aims to find the hyperparameters that minimize the validation loss when the training objective is trained to convergence. [^1] This notation assumes unique solutions for simplicity. See Vicol et al. (2022) for effects of non-uniqueness. Hyperparameter optimization can be done in sequential fashion, where a proposal $\boldsymbol{\lambda}_{n}$ depends on the sequence of prior values $\{\boldsymbol{\lambda}_{1},\boldsymbol{\lambda}_{2},\dots,\boldsymbol{\lambda}_{n-1}\}$ and their validation losses.
For instance, Bayesian Optimization (Mockus, 1998; Shahriari et al., 2015) builds a probabilistic model, such as a Gaussian Process, to map hyperparameters $\boldsymbol{\lambda}$ to the validation loss $\mathcal{L}_{V}(\boldsymbol{\lambda},\mathbf{w}^{*}(\boldsymbol{\lambda}))$. This approach iteratively selects the next hyperparameters to evaluate by optimizing an acquisition function that balances exploration and exploitation, thereby converging to the optimal hyperparameters $\boldsymbol{\lambda}^{*}$ that minimize the validation loss. In a limited budget setting, practitioners often employ a trial-and-error “manual" search, choosing hyperparameters based on prior knowledge or experience (Yang and Shami, 2020). In this paper, we assess the ability of large language models in this role, hypothesizing they will be effective as they have been trained on internet-scale data and have demonstrated emergent capabilities in new settings (Brown et al., 2020; OpenAI, 2023).
@@ 2.2 LLMs for Hyperparameter Optimization
We now describe our approach in more detail. We first address the case when the hyperparameters search space is known. We prompt an LLM by describing the problem and the search space. The LLM outputs a hyperparameter setting to evaluate. We then repeatedly alternate between two steps: (1) evaluating the current hyperparameter setting, and (2) prompting the LLM with the validation metric (e.g., loss or accuracy) to receive the next set of hyperparameters. This continues until our search budget is exhausted. We illustrate this process in Figure 1 and list the prompts that we use in Appendices B and D. Note that our prompts provide a high-level overview of the machine learning problem setup, the hyperparameters that can be tuned, and, if given, their associated search spaces.
@@
[User: <<Machine learning problem description>>. Provide a config in JSON format. Config:
@@
User: The validation loss was <<L>>. Provide the next config in the same format.
) x N]
@@
[User: <<Machine learning problem description>>. Here is what we have tried so far:
@@
Provide the next config in JSON format. Config:]
@@
More concretely, we consider two ways to prompt the model, following the popular “chat” interface Touvron et al. (2023a); OpenAI (2023) where users prompt the model with a dialogue consisting of “user” and “assistant” messages, depicted in Figure 1. The models we use generate tokens sampled according to an input temperature, terminating at a stop token. We prompt the model with the entire conversational history of messages, so the number of messages sent in the $n$-th prompt scales linearly with the number of steps. This is because every new step contains two additional messages: the previous LLM response and the validation metrics from executing a training run from the hyperparameters in the response. The number of tokens increases by the number of tokens representing the validation metrics and configuration. We denote this approach the “chat” prompt.
An alternative is to compress the previous search history into a single initial message. The single message contains the problem description and lists the configurations’ history and corresponding validation metrics. This approach represents a more compact state representation, especially if we use chain-of-thought reasoning (Wei et al., 2022) prompting strategies explored in our experiments. The inference cost on compressed messages is cheaper than in the prior approach. Empirically, we observe that the two approaches achieve similar performance. We denote this approach the “compressed” prompt.
When the search space is unknown, we can prompt an LLM to generate code for the model and optimizer, effectively providing a hyperparameter setting to evaluate. In this case, the environment executes the generated code and provides the resulting output to the LLM. We use the prompt in Appendix C. If the code fails to produce a validation accuracy, then we can re-prompt with the error message. Note that code generation can also be viewed as hyperparameter tuning, where the hyperparameter is the input string representing the code.
@@ 3 Related Work
Finding the optimal hyperparameter settings is crucial to achieving strong performance in machine learning (Bergstra and Bengio, 2012; Snoek et al., 2012). We refer readers to Feurer and Hutter (2019) and Bischl et al. (2023) for a general introduction to HPO. Initial research mainly explored model-free techniques such as grid and random search (Bergstra and Bengio, 2012). More advanced methods leverage multi-fidelity optimization – often by using that our optimization is an iterative process. For example, Hyperband (Li et al., 2017) and the Successive Halving (Jamieson and Talwalkar, 2016) introduced multi-armed bandit approaches to allow for the early stopping of less promising hyperparameter configurations. These methods are easy to parallelize and have shown promising results in practice, but are dependent on random processes and do not fully leverage the HPO structure.
Bayesian Optimization (BO) (Mockus, 1998; Hutter et al., 2011; Snoek et al., 2012; Bergstra et al., 2013; Shahriari et al., 2015) builds a surrogate model from past function evaluations to choose promising future candidates. BO further models uncertainty, leveraged by optimizing an acquisition function instead of the loss directly. Since the initial successes of BO in optimizing hyperparameters, numerous tools have been developed to optimize the pipeline for improved efficiency and adaptability (Bakshy et al., 2018; Kandasamy et al., 2020; Balandat et al., 2020; Lindauer et al., 2022). However, BO requires building an inductive bias in conditional probability, is heavily influenced by the choice of surrogate model parameters (Lindauer et al., 2019), and, importantly, faces scalability problems with an increase in the number of hyperparameters being tuned or the number of past function evaluations. Gradient or conditioning-based methods (Maclaurin et al., 2015; Franceschi et al., 2017; Lorraine and Duvenaud, 2018; MacKay et al., 2019; Bae and Grosse, 2020; Lorraine et al., 2020; Raghu et al., 2021; Bae et al., 2022) present a more scalable and efficient solution for HPO. Nevertheless, they are often challenging to implement, require the objective to be differentiable, and must be deployed in the same location as the underlying model, which makes them less appealing as general HPO solvers. A related approach is OptFormer (Chen et al., 2022), which finetunes Transformers on a large offline dataset to transfer learn a surrogate used for gradient-based HPO.
@@
Large language models (LLMs) have proven to be exceptionally valuable in a variety of practical domains (Hariri, 2023; Liu et al., 2023), showing surprising emergent abilities, including in-context learning and chain-of-thought reasoning (Brown et al., 2020; Wei et al., 2022; OpenAI, 2023). Although LLMs are known to occasionally give confident but incorrect answers (Ji et al., 2023), they are also shown to have reasoning capabilities, especially when explicitly guided to (Nye et al., 2021; Wei et al., 2022; Kojima et al., 2022). Recent studies have used LLMs for optimization (Yang et al., 2023), such as finding the best prompt for the downstream task or integrating LLMs into the general AutoML pipeline (Chen et al., 2023a; Hollmann et al., 2023; Lorraine et al., 2022; Zhang et al., 2023; Zheng et al., 2023). For example, Zheng et al. (2023) uses similar prompting strategies for neural architecture search (NAS) and demonstrates that LLMs can find competitive architectures on NAS benchmarks (Su et al., 2021).

### v2

@@ Abstract
This paper explores the use of foundational large language models (LLMs) in hyperparameter optimization (HPO). Hyperparameters are critical in determining the effectiveness of machine learning models, yet their optimization often relies on manual approaches in limited-budget settings. By prompting LLMs with dataset and model descriptions, we develop a methodology where LLMs suggest hyperparameter configurations, which are iteratively refined based on model performance. Our empirical evaluations on standard benchmarks reveal that within constrained search budgets, LLMs can match or outperform traditional HPO methods like Bayesian optimization across different models on standard benchmarks. Furthermore, we propose to treat the code specifying our model as a hyperparameter, which the LLM outputs and affords greater flexibility than existing HPO approaches.
@@ 1 Introduction
Hyperparameters, distinct from the parameters directly learned throughout training, significantly influence the inductive bias of a machine learning model, thus determining its capacity to generalize effectively. The choice of hyperparameters helps dictate the complexity of the model, the strength of regularization, the optimization strategy, the loss function, and more. Due to this importance, many methods have been proposed to find good hyperparameter configurations automatically. For example, black-box optimization methods such as random search (Bergstra & Bengio, 2012) or Bayesian optimization (Mockus, 1998; Shahriari et al., 2015) have been deployed for hyperparameter optimization (HPO). Beyond black-box methods, other works use problem structure, e.g., are compute-budget-aware (Lam et al., 2016), multi-task (Swersky et al., 2013), multi-fidelity (Klein et al., 2017; Li et al., 2017), online (Jaderberg et al., 2017), or multi-objective (Daulton et al., 2022). However, these methods (a) still rely on practitioners to design a search space, which includes selecting which parameters can be optimized and specifying bounds on these parameters, and (b) typically struggle in the initial search phase ($<2^{d}$ queries for $d$-dimensional hyperparameters). As such, tuning hyperparameters can be challenging for those with constrained budgets or without machine learning expertise, and methods that make HPO easier are desirable.
@@
This paper investigates the ability of large language models (LLMs) to optimize hyperparameters. LLMs are trained on internet-scale data and have demonstrated emergent capabilities in new settings (Brown et al., 2020; OpenAI, 2023). We prompt LLMs with an initial set of instructions—describing the specific dataset, model, and hyperparameters—to recommend hyperparameters to evaluate. We train the model based on the proposed hyperparameters, record the final metric (e.g., validation loss), and ask the LLM to suggest the next set of hyperparameters. Figure 1 illustrates this iterative process. We found that LLMs were effective at proposing hyperparameters during initial search phases, making them valuable both for limited-budget scenarios and as a complement to traditional approaches when larger search budgets are available.
@@
Our experiments demonstrated that LLMs can effectively optimize hyperparameters across a range of tasks, from toy problems to real-world applications. We first assessed LLMs’ ability to optimize 2D toy objectives, where they received loss $\mathcal{L}(\boldsymbol{x})$ at specific points $\boldsymbol{x}$. Across a range of objectives, we found that LLMs effectively minimized the loss, exploiting performant regions while exploring untested areas. Next, to analyze whether LLMs could optimize realistic HPO settings, we evaluated our approach on standard HPO benchmarks (Eggensperger et al., 2021), comparing it to common methods such as random search and Bayesian optimization. With small search budgets (e.g., $30$ evaluations), LLMs improved upon traditional hyperparameter tuning. We also assessed the importance of using chain-of-thought reasoning (Wei et al., 2022b) and demonstrated that LLMs performed well over longer horizons of up to $100$ evaluations. Additionally, LLMs were significantly faster than random search in tuning hyperparameters for both Vision Transformers and ResNets on the CIFAR-10 dataset, even when equipped with generic, architecture-agnostic prompts (“You are helping tune hyperparameters for a neural network.”).
@@
We further explored hyperparameter optimization using the unique flexibility that natural language enables. Rather than adhering to a fixed set of hyperparameter configurations, we prompted LLMs to produce training code (e.g., in PyTorch) to improve validation performance. This extension alleviates the need for human specification of the hyperparameters and their search spaces. The generated code can reduce the initial search for configuration spaces that are unlikely to succeed. With a limited search budget ($5$ evaluations), our results show that code generation performs better than random search and is comparable to LLM search with human-specified configurations. Finally, we conclude by discussing the potential of language models as general-purpose hyperparameter tuning assistants, limitations, and future work.
@@ 2 Method
LLMEnvironment Initial prompt: You are helping tune hyperparameters to minimize loss… this is our search space… formulate your output as JSON. Hyperparameter configuration Validation metric e.g., loss
Figure 1: LLMs for hyperparameter optimization. We prompt an LLM with the problem description and the search space. The LLM then outputs a set of hyperparameters to evaluate. The environment, e.g., practitioner or automatic script, executes a training run with the hyperparameter setting, and then a validation metric is used to prompt the language model again.
@@ 2.1 Background
$\boldsymbol{\lambda}^{\star}$ $=\operatorname*{arg\,min}_{\boldsymbol{\lambda}}\mathcal{L}_{V}^{*}\left(\boldsymbol{\lambda}\right)=\operatorname*{arg\,min}_{\boldsymbol{\lambda}}\mathcal{L}_{V}\left(\boldsymbol{\lambda},\mathbf{w}^{\star}\left(\boldsymbol{\lambda}\right)\right)$
s.t. $\mathbf{w}^{\star}\left(\boldsymbol{\lambda}\right)=\operatorname*{arg\,min}_{\mathbf{w}}\mathcal{L}_{T}\left(\boldsymbol{\lambda},\mathbf{w}\right),$
where $\mathcal{L}_{T}$ and $\mathcal{L}_{V}$ are training and validation objectives, and $\boldsymbol{\lambda}$ and $\mathbf{w}$ are hyperparameters and the model parameters, respectively. The objective aims to find the hyperparameters that minimize the validation loss when the training objective is trained to convergence. [^1] This assumes unique solutions for simplicity. See Vicol et al. (2022) for effects of non-uniqueness.
Hyperparameter optimization can often be performed sequentially, where a proposal $\boldsymbol{\lambda}_{n}$ depends on the sequence of prior values $\{\boldsymbol{\lambda}_{1},\boldsymbol{\lambda}_{2},\dots,\boldsymbol{\lambda}_{n-1}\}$ and their validation losses. For example, Bayesian optimization (Mockus, 1998; Shahriari et al., 2015) builds a probabilistic model, such as a Gaussian process, to map hyperparameters $\boldsymbol{\lambda}$ to the validation loss $\mathcal{L}_{V}(\boldsymbol{\lambda},\mathbf{w}^{*}(\boldsymbol{\lambda}))$. This approach iteratively selects the next hyperparameters to evaluate by optimizing an acquisition function that balances exploration and exploitation, thus converging to the optimal hyperparameters $\boldsymbol{\lambda}^{*}$ that minimize the validation loss. In a limited budget setting, practitioners often employ a trial-and-error “manual” search, choosing hyperparameters based on prior knowledge or experience (Yang & Shami, 2020). In this paper, we assess the ability of large language models in this role, hypothesizing they will be effective because they have been trained on internet-scale data and demonstrated emergent capabilities in new settings (Brown et al., 2020; Wei et al., 2022a).
@@ 2.2 LLMs for Hyperparameter Optimization
We now describe our approach in more detail. When the hyperparameter search space is known, we prompt an LLM with a description of the problem and the search space. The LLM outputs a hyperparameter setting to evaluate. We then repeatedly alternate between two steps: (1) evaluating the current hyperparameter setting and (2) prompting the LLM with the validation metric (e.g., loss or accuracy) to receive the next set of hyperparameters. This continues until our search budget is exhausted. We illustrate this process in Figure 1 and list our prompts in Appendices B, D, and E. The prompts provide a high-level overview of the machine learning problem setup, the hyperparameters to be tuned, and, if given, their search spaces. This generality enables our approach to be applied to many domains. Example code can be found at https://github.com/michaelrzhang/LLM-HyperOpt.
@@
User: <<Machine learning problem description>>. Provide a config in JSON format. Config:
@@
User: The validation loss was <<L>>. Provide the next config in the same format.) x N
@@
User: <<Machine learning problem description>>. Here is what we have tried so far:
@@
Provide the next config in JSON format. Config:
@@
More concretely, we consider two ways to prompt the model, following the popular “chat” interface (Touvron et al., 2023a; OpenAI, 2023) where users prompt the model with a dialogue consisting of “user” and “assistant” messages, depicted in Figure 2. The models we use generate tokens sampled according to an input temperature, terminating at a stop token. In the first approach, we prompt the model with the entire conversational history of messages, so the number of messages sent scales linearly with the number of steps. This happens because every new step contains two additional messages: the previous LLM response and the validation metrics from executing a training run from the hyperparameters in the response. The inference cost is linear (with a constant attention window) or quadratic in the number of tokens for a Transformer (Brown et al., 2020). For every new iteration, the number of tokens increases by the tokens representing the validation metrics and configuration. We denote this approach the “chat” prompt.
An alternative prompting strategy is to compress the search history into a single initial message that contains the problem description and lists the hyperparameter configuration history and corresponding validation metrics. This approach provides a more compact state representation, especially if we use chain-of-thought prompting (Wei et al., 2022b) to elicit reasoning, as in some experiments. The inference cost on compressed messages is cheaper than in the prior approach. Empirically, we observe that the two approaches achieve similar performance. We denote this approach as the “compressed” prompt and provide an implementation and example usage in our supplementary material.
When the search space is unknown, we can prompt an LLM to generate code for the model and optimizer, effectively providing a hyperparameter setting to evaluate. In this case, the environment, e.g., the practitioner, executes the generated code and provides the resulting output to the LLM. We use the prompt in Appendix D that asks the LLM to write code in PyTorch code that defines a model and optimizer. If the code fails to produce a valid output, we can re-prompt with the error message. Code generation can also be viewed as hyperparameter tuning, where the hyperparameter $\boldsymbol{\lambda}$ is the input string representing the code. This extension alleviates the need for human specification of the hyperparameters to tune and their search spaces.
@@ 3 Related Work
Finding the optimal hyperparameter settings is crucial to achieving strong performance in machine learning (Bergstra & Bengio, 2012; Snoek et al., 2012). We refer readers to Feurer & Hutter (2019) and Bischl et al. (2023) for a general introduction to HPO. Initial research mainly explored model-free techniques such as grid and random search (Bergstra & Bengio, 2012). More advanced methods leverage multi-fidelity optimization – often by using that our optimization is an iterative process. For example, Hyperband (Li et al., 2017) and the Successive Halving (Jamieson & Talwalkar, 2016) introduced multi-armed bandit approaches to allow for the early stopping of less promising hyperparameter configurations. These methods are easy to parallelize and have shown promising results in practice, but they are dependent on random processes and do not fully leverage the HPO structure.
Bayesian Optimization (BO) (Mockus, 1998; Hutter et al., 2011; Snoek et al., 2012; Bergstra et al., 2013; Shahriari et al., 2015) builds a surrogate model from past function evaluations to choose promising future candidates. Bayesian optimization further models uncertainty, leveraged by optimizing an acquisition function instead of the loss directly. Since the initial successes of BO in optimizing hyperparameters, numerous tools have been developed to optimize the pipeline for improved efficiency and adaptability (Bakshy et al., 2018; Kandasamy et al., 2020; Balandat et al., 2020; Lindauer et al., 2022). However, Bayesian optimization is heavily influenced by the choice of surrogate model parameters (Lindauer et al., 2019), and faces scalability problems with an increased hyperparameter number being tuned or the number of past function evaluations. We show in our experiments that LLMs can exceed the performance of BO in the initial phase of tuning hyperparameters and the potential benefit of adopting a hybrid strategy of two algorithms.
Gradient or conditioning-based methods (Maclaurin et al., 2015; Franceschi et al., 2017; Lorraine & Duvenaud, 2018; MacKay et al., 2019; Bae & Grosse, 2020; Lorraine et al., 2020; Raghu et al., 2021; Bae et al., 2022; Mehta et al., 2024; Lorraine, 2024) are more scalable and efficient HPO methods. Nevertheless, they are often challenging to implement, require a differentiable objective, and must be deployed in the same location as the underlying model, making them less appealing general HPO solvers. A related method is OptFormer (Chen et al., 2022), which finetunes transformers on a large offline dataset to transfer learn a surrogate for gradient-based HPO.
@@
Large language models (LLMs) have proven to be valuable in a variety of practical domains (Hariri, 2023; Liu et al., 2023), showing surprising emergent abilities, including in-context learning and chain-of-thought reasoning (Brown et al., 2020; Wei et al., 2022b; OpenAI, 2023). Although LLMs are known to occasionally give confident but incorrect answers (Ji et al., 2023), they are also shown to have reasoning capabilities, especially when explicitly guided to (Nye et al., 2021; Wei et al., 2022b; Kojima et al., 2022). Recent studies have used LLMs for optimization (Yang et al., 2024), such as finding the best downstream task prompt. They conducted experiments on linear regression, Traveling Salesman, and prompt optimization. In contrast, our work focuses on using LLMs to make decisions during hyperparameter optimization for machine learning tasks, including classical ML algorithms and training deep neural networks.
Others have integrated LLMs into the general AutoML pipeline (Chen et al., 2023; Hollmann et al., 2023; Lorraine et al., 2022; Zhang et al., 2023; Zheng et al., 2023). For example, Zheng et al. (2023) demonstrate that LLMs can find competitive architectures on neural architecture search benchmarks (Su et al., 2021).